OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form

OpenStudy (anonymous):

y-2=-(x+4) y-2=-x-4 y+x=-2

OpenStudy (amistre64):

parallels are simple enough, they differ by a constant; so cut off that constant and re calibrate with the given point

OpenStudy (amistre64):

given (–4, 2) and: y = –x + 6 ; remove the constant y = –x ; calibrate with the given point 2 = -(-4) + C 2 = 4 + C ; C = -2 parallel line equation is: y = –x - 2

OpenStudy (anonymous):

thanxs for all the help